If x=5 in the expression 5x+2y=7 what the value of y?
Answer by elima(1433)

If x=5 in the expression 5x+2y=7 what the value of y?
Since you are given what x =, then plug that into the equation;
5(5)+2y=7
25+2y=7
2y=7-25
2y=-18
y=-9
:)